# Enabling the Write-Back Page Cache with Strong Consistency in Distributed Userspace File Systems

### v2

@@ Abstract.
Cloud platforms host thousands of tenants that demand POSIX semantics, high throughput, and rapid evolution from their storage layer. Kernel-native distributed file systems supply raw speed, but their privileged code base couples every release to the kernel, widens the blast radius of crashes, and slows innovation. FUSE-based distributed file systems flip those trade-offs: they run in user space for fast deployment and strong fault isolation, yet the FUSE interface disables the kernel’s write-back page cache whenever strong consistency is required. Practitioners must therefore choose between (i) weak consistency with fast write-back caching or (ii) strong consistency with slow write-through I/O, an limitation that has kept FUSE distributed file systems out of write-intensive cloud workloads.
To this end, We present DistFUSE, the first distributed FUSE file system that delivers write-back kernel caching and strong consistency. DistFUSE achieves this by offloading userspace consistency control to the kernel driver, allowing coordinated access to the kernel’s page cache across nodes. This design eliminates blind local cache updates and ensures cluster-wide strong consistency without compromising performance. In our evaluation, DistFUSE achieves up to 68.0% higher throughput and 40.4% lower latency than the existing write-through design of FUSE-based distributed file system.
FUSE, write-back cache, consistency, distributed file systems
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineering File systems managementccs: Software and its engineering Consistencyccs: Computer systems organization Cloud computingccs: Software and its engineering Distributed systems organizing principlesccs: Information systems Distributed storage
@@ 1. Introduction
Table 1. Comparison between DistFUSE and existing solutions.
@@
Dimension DistFUSE Ceph-fuse GlusterFS Lustre GPFS
@@
However, a fundamental challenge persists: FUSE-based distributed file systems struggle to provide both high performance and strong consistency (i.e., allow distributed clients to read the most recent written value). The kernel’s page cache—a key driver of efficiency in local file systems—remains largely incompatible with distributed environments. The write-back page cache enables applications to return from I/O operations immediately after buffering data in the kernel, removing the expensive userspace synchronization of FUSE from the critical path. Yet, in distributed settings, this risks data races and inconsistency, as nodes may operate on stale or uncoordinated cached copies. For example, if node A writes data to its local page cache (happens in the kernel) without immediately synchronizing with other nodes (happens in the userspace), node B might read stale data or make conflicting updates, leading to data corruption. Using the write-through mode of kernel page cache mitigates the consistency issue at cost of low performance (Figure 2). Therefore, prior solutions either sacrifice strong consistency for performance (e.g., enabling a write-back cache without synchronization)(noa, 2011, 2025g) or default to write-through caching(Weil et al., 2006), which synchronizes every write to userspace (Table 1). In write-through mode, each operation must complete synchronization before returning to the application, incurring high latency. This trade-off limits practicality in consistency-sensitive cloud workloads, such as databases or real-time analytics.
@@
We present DistFUSE, a FUSE-based distributed file system that bridges the gap between high performance and strong consistency. DistFUSE enables write-back page caching while enforcing strong consistency across nodes, eliminating the need to compromise between performance and correctness. The core insight of DistFUSE is offloading userspace consistency control (e.g., lease management) to the kernel FUSE driver. By embedding distributed metadata directly into the FUSE driver, DistFUSE coordinates access permissions within the kernel. This design resolves the kernel space deadlock issue in current systems (discussed in §3.2). This allows nodes to gain cluster-wide access within the kernel during I/O operations, ensuring that only authorized nodes read or modify cached data. For instance, when a write lease is revoked, the kernel proactively invalidates the local page cache and flushes dirty pages. This design retains FUSE’s userspace flexibility while leveraging the kernel’s efficiency, avoiding costly round trips for common-case operations.
@@
Our evaluation demonstrates that DistFUSE consistently improves I/O throughput and latencies compared to a baseline, which represents the architecture of systems such as GlusterFS(noa, 2025g), an open-source distributed file system based on FUSE. Results demonstrate a 1.75x throughput improvement at best, outperforming write-through approaches. DistFUSE also scales linearly with node count, maintaining consistency even at 25% contention.
@@
  • The observation that leveraging kernel’s page cache fundamentally breaks strong consistency in FUSE-based distributed file systems.
@@
  • The design and implementation of DistFUSE, the first FUSE-based system to support write-back caching with strong consistency.
The rest of this paper is organized as follows: §2 provides background on FUSE and FUSE-based distributed file systems; §3 details consistency and locking challenges; §4 and §5 present DistFUSE’s design and implementation, respectively; §6 evaluates its efficacy; §7 discusses related work; and §8 concludes the paper.
@@ 2.2. FUSE Architecture Overview
FUSE is a framework that allows developers to build userspace filesystems exportable to Linux kernel. As illustrated in Fig 1, it comprises three key components: the FUSE kernel driver, the userspace library (libfuse (noa, 2025i)), and a custom FUSE daemon. The kernel driver registers the file system with the VFS, allowing applications to interact with it via standard POSIX interfaces (e.g., open, read, write) exposed by the GNU C Library (glibc). When an application issues a syscall, glibc invokes the corresponding VFS function, which routes the request to the FUSE driver. The driver translates the syscall into a FUSE protocol message (e.g., FUSE_READ) and forwards it to userspace via a special device called /dev/fuse. To summarize, while FUSE is a userspace-based file system, when applications interact with it, the I/O request gets routed first to the kernel then back to the file system in userspace.
@@ 2.4. FUSE Performance Breakdown
We quantify the performance benefits of kernel’s write-back page cache in Fig 2. To break down the overhead of FUSE protocol itself, we implement a NullFS which simply returns zero in any file system operations in the userspace. Note that none of the requests actually go to the disk. The stacked bars (from the bottom up) present the life cycles of write requests chronologically. [^1] The workload is 1M random 4KB writes over 1K files of 16MB after warm-up. First, the application issues a write system call, which triggers a mode switch into the kernel, and the VFS layer dispatches the request to the FUSE driver. In both cache modes, the content within the write request will be copied to the kernel’s page cache. If the write-back cache is enabled, FUSE then returns the number of written bytes or an error to the application, with an average total latency of 4.7 $\mu$s (the sum of the left bar).
In contrast, in the write-through mode, after writing the page cache, FUSE continues to pass the write request to the userspace daemon. In particular, the FUSE driver as in Fig 1 first places the request in one of the internal queues and next wakes up the daemon thread that waits on the queues (7.2 $\mu$s). Subsequently, the daemon thread dequeues the request and copies it to the userspace buffer (2.7 $\mu$s). After switching to the userspace, the daemon invokes the userspace write function to handle the requests (2.5 $\mu$s). Finally, the response returns to the kernel driver in the same way, i.e., the daemon copies the reply to the kernel buffer (0.7 $\mu$s) and notifies the FUSE driver thread of the incoming result (6.1 $\mu$s).
From Fig 2, the overhead of the extra request round trip in FUSE write-through cache is on average 19.2 $\mu$s, while write-back cache mode incurs only 4.7 $\mu$s.
@@ 2.5. Achieving Strong Consistency
A common approach used in distributed file systems is the lease mechanism (Gray and Cheriton, 1989; Schmuck and Haskin, 2002; Schwan et al., 2003; Kronenberg et al., 1986; Mohindra and Devarakonda, 1994). Compared to other synchronization mechanism, e.g., distributed locks, leases achieve lower synchronization overhead as the lease-holding node can safely perform operations locally during a lease period without communication with the lease coordinator. Once the lease expires or is revoked, the node must either renew it or cease operations on the leased data until it obtains a new lease.
In FUSE-based distributed file systems, implementing leases is particularly challenging due to the separation between kernel-resident page caches and userspace consistency logic. For instance, a node holding a write lease may buffer updates in its kernel page cache, leaving other nodes unaware of pending changes until a flush occurs. Prior systems like Ceph-FUSE resort to write-through caching to synchronize every update with userspace, while Gluster sacrifices consistency for performance. We elaborate on these challenges in §3 and explain how DistFUSE addresses these limitations by embedding lease management directly into the kernel driver, enabling write-back caching without compromising consistency in §4.
@@ 3.2. Deadlock Caused by Reversed Lock Order With Write Through
Figure 3(c) presents an example. Suppose that node N has held a write lease for file f from the beginning. When writing f, it first need to grab the inode lock in the kernel (step 2). Once the write request arrives in the userspace, it has to obtain another lock of the userspace lease to read its content (step 4). This kernel-userspace locking order would conflict with the sequence required during a lease revocation request. In particular, if another node wants to write the same file f (step I) and tries to acquire the distributed lease of it (step II), an revoke request for file f’s lease may be sent to node $N$ (step III). At this time, the revoking request will grab locks in the reversed order: first locking the lease (step IV) and then the inode in the kernel (step V), which leads to dead locks.
@@ 4. Design
Figure 4. The overview of DistFUSE
DistFUSE is a FUSE-based distributed file system that enables write-back page caching while maintaining strong consistency. The system bridges the gap between high performance and strong consistency by offloading distributed concurrency control to the kernel driver, allowing the kernel’s page cache to be safely utilized across distributed nodes. Table 1 highlights DistFUSE’s differences from existing FUSE-based and kernel-native solutions. Figure 4 illustrates DistFUSE’s overall architecture, which consists of three major components:
@@
These components run on different physical servers and communicate with each other via RPCs. To use DistFUSE, an application simply mounts a DFS client to a local directory (e.g., /mnt/fuse) and accesses files as if using a local filesystem. Applications across all nodes see a unified and consistent view of the filesystem through the standard POSIX interface.
@@ 4.1. DFS Client
Each DFS client functions as a FUSE-based file system mounted to a local directory, collectively providing a unified namespace across all nodes. Applications interact with DistFUSE using standard POSIX interfaces. The key innovation in DistFUSE is offloading userspace distributed lease management to the kernel driver, enabling the system to validate lease status directly within the kernel during I/O operations.
This lease offloading mechanism is what enables DistFUSE to use write-back page caching while maintaining strong consistency across the distributed system. Figure 5 illustrates the internal components of a DFS client. Importantly, DistFUSE confines all kernel modifications to the FUSE kernel driver, leaving other kernel components (e.g., VFS and page cache) unchanged. Users simply need to load the DistFUSE kernel driver and install our customized libfuse.
4.1.1. Why DistFUSE chooses leases
Distributed file systems typically enforce cache consistency with callback locks (Howard et al., 1988), leases (Gray and Cheriton, 1989), version vectors (Shepler et al., 2003), or optimistic validation (Kung and Robinson, 1981). We adopt leases because they already balance low-latency reads with bounded staleness and are widely deployed in production DFSs (Ghemawat et al., 2003; Shvachko et al., 2010; Calder et al., 2011; Corbett et al., 2012; Rumble et al., 2012). The lease protocol itself is therefore not a contribution of this work; its role is to provide a familiar yard-stick against which to evaluate DistFUSE’s kernel-offload mechanism. DistFUSE’s design also has the potential to utilize other consistency protocols mentioned above.
@@ 4.1.2. Lease Offloading
DistFUSE’s core innovation is embedding distributed lease information directly in the kernel’s FUSE driver. In this section, we show that how this design allows the system to coordinate page cache access across nodes without sacrificing performance.
DistFUSE adds a distributed lease field to the FUSE inode data structure, representing the current node’s access permission for that inode. This lease has three possible states: shared read, exclusive write, or null. The lease initializes to null when an application opens a file, indicating no initial read or write access.
For each inode operation requiring shared read or exclusive write, the DistFUSE driver checks the lease type after acquiring the kernel inode lock. If it finds that the node is not holding the appropriate lease (null or weaker than required), the driver requests the userspace daemon to update the lease. Conversely, when the DistFUSE daemon receives a revocation from the lease manager, it instructs the kernel driver to release the lease.
To manage these offloaded leases, DistFUSE introduces two new FUSE operations:
@@
Called when the driver determines its local lease does not provide sufficient permission. The driver packages the inode number and desired permission (read or write) into a request sent to the userspace daemon through /dev/fuse. The daemon then acquires the distributed lease from the lease manager via RPC. Upon success, the driver updates the local lease and proceeds with the application’s operation. The application blocks on a read or write until the FUSE driver finish the operation.
@@
To better demonstrate the lease offloading mechanism in DistFUSE, we show the workflow of a write request from an application running on DFS client A. First, the application issues a write syscall with glibc, which will be routed to the FUSE kernel driver by VFS. Next, the FUSE driver locks the lease of the inode (serialize concurrent revocation) and checks the lease status. If the lease has the write permission, then the FUSE driver proceeds with the normal write-back operation, i.e., acquiring the inode lock and updating the page cache. However, if the lease staus is either null or shared read, the FUSE driver will send a fuse_acquire_dist_lease() operation to the userspace daemon through /dev/fuse. The kernel thread falls into interruptible sleep while waiting for the result. Once the userspace daemon acquires a write lease from the manager, it wakes up the FUSE driver thread. Lastly, the driver updates the lease type to write, updates the kernel page cache, and returns the control to the application. This solves the challenge mentioned in §3.1.
@@
We next discuss how a DFS client handles lease revocations. Suppose that during a write request, the userspace daemon of client A receives a lease revocation request for the same file. To properly revoke the lease, client A must avoid new I/Os to this file and the deadlock issue mentioned in §3.2. So, the daemon triggers a fuse_release_dist_lease() notification to the FUSE driver, which will first lock the lease (to prevent new I/Os from reading the status) and then lock the inode (to flush its cached pages). By following the same locking order as normal read and write requests, DistFUSE resolves the deadlock challenge in §3.2. Subsequent I/Os to the file will find the lease in the null status and trigger the above lease acquisition.
4.1.3. Userspace Page Cache

### v3

# DFUSE: Strongly Consistent Write-Back Kernel Caching for Distributed Userspace File Systems

@@ Abstract.
Cloud platforms host thousands of tenants that demand POSIX semantics, high throughput, and rapid evolution from their storage layer. Kernel-native distributed file systems supply raw speed, but their privileged code base couples every release to the kernel, widens the blast radius of crashes, and slows innovation. FUSE-based distributed file systems flip those trade-offs: they run in user space for fast deployment and strong fault isolation, yet the FUSE interface disables the kernel’s write-back page cache whenever strong consistency is required. Practitioners must therefore choose between (i) weak consistency with fast write-back caching or (ii) strong consistency with slow write-through I/O, a limitation that has kept FUSE distributed file systems out of write-intensive cloud workloads.
To this end, we present DFUSE, the first distributed FUSE file system that delivers write-back kernel caching and strong consistency. DFUSE achieves this by offloading userspace consistency control to the kernel driver, allowing coordinated access to the kernel’s page cache across nodes. This design eliminates blind local cache updates and ensures cluster-wide strong consistency without compromising performance. In our evaluation, DFUSE achieves up to 68.0% higher throughput and 40.4% lower latency than the existing write-through design of FUSE-based distributed file systems.
FUSE, cache, consistency, distributed file systems
journalyear: 2025copyright: acmlicensedconference: ACM Symposium on Cloud Computing; November 19–21, 2025; Online, USAbooktitle: ACM Symposium on Cloud Computing (SoCC ’25), November 19–21, 2025, Online, USAdoi: 10.1145/3772052.3772208isbn: 979-8-4007-2276-9/25/11ccs: Software and its engineering File systems managementccs: Software and its engineering Consistencyccs: Information systems Distributed storageccs: Computer systems organization Cloud computingccs: General and reference Performance
@@ 1. Introduction
Table 1. Comparison between DFUSE and existing solutions.
@@
Dimension DFUSE Ceph-fuse GlusterFS Lustre GPFS
@@
However, a fundamental challenge persists: FUSE-based distributed file systems struggle to provide both high performance and strong consistency (i.e., allow distributed clients to read the most recent written value). The kernel’s page cache—a key driver of efficiency in local file systems—remains largely incompatible with distributed environments. The write-back page cache enables applications to return from I/O operations immediately after buffering data in the kernel, removing the expensive userspace synchronization of FUSE from the critical path. Yet, in distributed settings, this risks data races and inconsistency, as nodes may operate on stale or uncoordinated cached copies. For example, if node A writes data to its local page cache (happens in the kernel) without immediately synchronizing with other nodes (happens in the userspace), node B might read stale data or make conflicting updates, leading to data corruption. Using the write-through mode of kernel page cache mitigates the consistency issue at the cost of low performance (Figure 2). Therefore, prior solutions either sacrifice strong consistency for performance (e.g., enabling a write-back cache without synchronization)(noa, 2011, 2025g) or default to write-through caching(Weil et al., 2006), which synchronizes every write to userspace (Table 1). In write-through mode, each operation must complete synchronization before returning to the application, incurring high latency. This trade-off limits practicality in consistency-sensitive cloud workloads, such as databases or real-time analytics.
@@
We present DFUSE, a FUSE-based distributed file system that bridges the gap between high performance and strong consistency. DFUSE enables write-back page caching while enforcing strong consistency across nodes, eliminating the need to compromise between performance and correctness. The core insight of DFUSE is offloading userspace consistency control (e.g., lease management) to the kernel FUSE driver. By embedding distributed metadata directly into the FUSE driver, DFUSE coordinates access permissions within the kernel. This design resolves the kernel space deadlock issue in current systems (discussed in §3.2). This allows nodes to gain cluster-wide access within the kernel during I/O operations, ensuring that only authorized nodes read or modify cached data. For instance, when a write lease is revoked, the kernel proactively invalidates the local page cache and flushes dirty pages. This design retains FUSE’s userspace flexibility while leveraging the kernel’s efficiency, avoiding costly round trips for common-case operations.
@@
Our evaluation demonstrates that DFUSE consistently improves I/O throughput and latencies compared to a baseline, which represents the architecture of systems such as GlusterFS(noa, 2025g), an open-source distributed file system based on FUSE. Results demonstrate a 1.75x throughput improvement at best, outperforming write-through approaches. DFUSE also scales linearly with node count, maintaining consistency even at 25% contention.
@@
  • An analysis showing that leveraging kernel’s page cache fundamentally breaks strong consistency in FUSE-based distributed file systems.
@@
  • The design and implementation of DFUSE, the first FUSE-based system to support write-back caching with strong consistency.
The rest of this paper is organized as follows: §2 provides background on FUSE and FUSE-based distributed file systems; §3 details consistency and locking challenges; §4 and §5 present DFUSE’s design and implementation, respectively; §6 evaluates its efficacy; §7 discusses related work; and §8 concludes the paper.
@@ 2.2. FUSE Architecture Overview
FUSE is a framework that allows developers to build userspace file systems exportable to the Linux kernel. As illustrated in Fig 1, it comprises three key components: the FUSE kernel driver, the userspace library (libfuse (noa, 2025i)), and a custom FUSE daemon. The kernel driver registers the file system with the VFS, allowing applications to interact with it via standard POSIX interfaces (e.g., open, read, write) exposed by the GNU C Library (glibc). When an application issues a syscall, glibc invokes the corresponding VFS function, which routes the request to the FUSE driver. The driver translates the syscall into a FUSE protocol message (e.g., FUSE_READ) and forwards it to userspace via a special device called /dev/fuse. To summarize, while FUSE is a userspace-based file system, when applications interact with it, the I/O request gets routed first to the kernel then back to the file system in userspace.
@@ 2.4. FUSE Performance Breakdown
[…]
@@ 2.5. Achieving Strong Consistency
A common approach used in distributed file systems is the lease mechanism (Gray and Cheriton, 1989; Schmuck and Haskin, 2002; Schwan et al., 2003; Kronenberg et al., 1986; Mohindra and Devarakonda, 1994). Compared to other synchronization mechanisms, e.g., distributed locks, leases achieve lower synchronization overhead as the lease-holding node can safely perform operations locally during a lease period without communication with the lease coordinator. Once the lease expires or is revoked, the node must either renew it or cease operations on the leased data until it obtains a new lease.
In FUSE-based distributed file systems, implementing leases is particularly challenging due to the separation between kernel-resident page caches and userspace consistency logic. For instance, a node holding a write lease may buffer updates in its kernel page cache, leaving other nodes unaware of pending changes until a flush occurs. Prior systems like Ceph-FUSE resort to write-through caching to synchronize every update with userspace, while Gluster sacrifices consistency for performance. We elaborate on these challenges in §3 and explain how DFUSE addresses these limitations by embedding lease management directly into the kernel driver, enabling write-back caching without compromising consistency in §4.
@@ 3.2. Deadlock Caused by Reversed Lock Order With Write Through
Figure 3(c) presents an example. Suppose that node N has held a write lease for file f from the beginning. When writing f, it first needs to grab the inode lock in the kernel (step 2). Once the write request arrives in the userspace, it has to obtain another lock of the userspace lease to read its content (step 4). This kernel-userspace locking order would conflict with the sequence required during a lease revocation request. In particular, if another node wants to write the same file f (step I) and tries to acquire the distributed lease of it (step II), a revoke request for file f’s lease may be sent to node $N$ (step III). At this time, the revoking request will grab locks in the reversed order: first locking the lease (step IV) and then the inode in the kernel (step V), which leads to deadlocks.
@@ 4. Design
Figure 4. The overview of DFUSE
DFUSE is a FUSE-based distributed file system that enables write-back page caching while maintaining strong consistency. The system bridges the gap between high performance and strong consistency by offloading distributed concurrency control to the kernel driver, allowing the kernel’s page cache to be safely utilized across distributed nodes. Table 1 highlights DFUSE’s differences from existing FUSE-based and kernel-native solutions. Figure 4 illustrates DFUSE’s overall architecture, which consists of three major components:
@@
These components run on different physical servers and communicate with each other via RPCs. To use DFUSE, an application simply mounts a DFS client to a local directory (e.g., /mnt/fuse) and accesses files as if using a local file system. Applications across all nodes see a unified and consistent view of the file system through the standard POSIX interface.
@@ 4.1. DFS Client
Each DFS client functions as a FUSE-based file system mounted to a local directory, collectively providing a unified namespace across all nodes. Applications interact with DFUSE using standard POSIX interfaces. The key innovation in DFUSE is offloading userspace distributed lease management to the kernel driver, enabling the system to validate lease status directly within the kernel during I/O operations.
This lease offloading mechanism is what enables DFUSE to use write-back page caching while maintaining strong consistency across the distributed system. Figure 5 illustrates the internal components of a DFS client. Importantly, DFUSE confines all kernel modifications to the FUSE kernel driver, leaving other kernel components (e.g., VFS and page cache) unchanged. Users simply need to load the DFUSE kernel driver and install our customized libfuse.
4.1.1. Why DFUSE chooses leases
Distributed file systems typically enforce cache consistency with callback locks (Howard et al., 1988), leases (Gray and Cheriton, 1989), version vectors (Shepler et al., 2003), or optimistic validation (Kung and Robinson, 1981). We adopt leases because they already balance low-latency reads with bounded staleness and are widely deployed in production DFSs (Ghemawat et al., 2003; Shvachko et al., 2010; Calder et al., 2011; Corbett et al., 2012; Rumble et al., 2012). The lease protocol itself is therefore not a contribution of this work; its role is to provide a familiar yard-stick against which to evaluate DFUSE’s kernel-offload mechanism. DFUSE’s design also has the potential to utilize other consistency protocols mentioned above.
@@ 4.1.2. Lease Offloading
DFUSE’s core innovation is embedding distributed lease information directly in the kernel’s FUSE driver. In this section, we show how this design allows the system to coordinate page cache access across nodes without sacrificing performance.
DFUSE adds a distributed lease field to the FUSE inode data structure, representing the current node’s access permission for that inode. This lease has three possible states: shared read, exclusive write, or null. The lease initializes to null when an application opens a file, indicating no initial read or write access.
For each inode operation requiring shared read or exclusive write, the DFUSE driver checks the lease type after acquiring the kernel inode lock. If it finds that the node is not holding the appropriate lease (null or weaker than required), the driver requests the userspace daemon to update the lease. Conversely, when the DFUSE daemon receives a revocation from the lease manager, it instructs the kernel driver to release the lease.
To manage these offloaded leases, DFUSE introduces two new FUSE operations:
@@
Called when the driver determines its local lease does not provide sufficient permission. The driver packages the inode number and desired permission (read or write) into a request sent to the userspace daemon through /dev/fuse. The daemon then acquires the distributed lease from the lease manager via RPC. Upon success, the driver updates the local lease and proceeds with the application’s operation. The application blocks on a read or write until the FUSE driver finishes the operation.
@@
To better demonstrate the lease offloading mechanism in DFUSE, we show the workflow of a write request from an application running on DFS client A. First, the application issues a write syscall with glibc, which will be routed to the FUSE kernel driver by VFS. Next, the FUSE driver locks the lease of the inode (serialize concurrent revocation) and checks the lease status. If the lease has the write permission, then the FUSE driver proceeds with the normal write-back operation, i.e., acquiring the inode lock and updating the page cache. However, if the lease status is either null or shared read, the FUSE driver will send a fuse_acquire_dist_lease() operation to the userspace daemon through /dev/fuse. The kernel thread falls into interruptible sleep while waiting for the result. Once the userspace daemon acquires a write lease from the manager, it wakes up the FUSE driver thread. Lastly, the driver updates the lease type to write, updates the kernel page cache, and returns the control to the application. This solves the challenge mentioned in §3.1.
@@
We next discuss how a DFS client handles lease revocations. Suppose that during a write request, the userspace daemon of client A receives a lease revocation request for the same file. To properly revoke the lease, client A must avoid new I/Os to this file and the deadlock issue mentioned in §3.2. So, the daemon triggers a fuse_release_dist_lease() notification to the FUSE driver, which will first lock the lease (to prevent new I/Os from reading the status) and then lock the inode (to flush its cached pages). By following the same locking order as normal read and write requests, DFUSE resolves the deadlock challenge in §3.2. Subsequent I/Os to the file will find the lease in the null status and trigger the above lease acquisition.
4.1.3. Userspace Buffer Cache